# Unsupervised Machine Learning for Classifying CHIME Fast Radio Bursts and Investigating Empirical Relations

### v1

@@ 5 Conclusion and Future Prospect
Machine learning is a powerful tool to classify FRBs. In this paper, we applied unsupervised learning methods, including dimensionality reduction and clustering algorithms, to differentiate between repeaters and non-repeaters in the first CHIME/FRB catalog (CHIME/FRB Collaboration et al., 2021) and the CHIME/FRB Collaboration (2023) catalog (CHIME/FRB Collaboration et al., 2023). We extracted 16 parameters from the FRBs to serve as input features for unsupervised learning, ensuring that the information from the FRBs was sufficiently comprehensive. Ultimately, we successfully identified several candidate repeaters among the non-repeaters. Using the UMAP+k-means method, we identified 269 non-repeaters as repeater candidates, with an estimated repeating rate of 61.7%. With the UMAP+HDBSCAN method, 141 non-repeaters were identified as repeater candidates, yielding a repeating rate of 37.9%. Additionally, we found that FRBs in repeater clusters and non-repeater clusters exhibit different distributions across several features, suggesting that repeaters and non-repeaters may belong to distinct categories.
@@
We used six previously classified as non-repeaters but actually confirmed as repeaters FRB sources to evaluate the predictive capability of our model. The UMAP + k-means method successfully predicted five of these sources, while the UMAP + HDNSCAN method successfully predicted four. The only exception was FRB 20180910A, which could not be predicted. The reason for this is that many of its characteristics, such as frequency bandwidth, spectral index, and spectral running, closely resemble those of non-repeaters, making it distinctly different from typical repeaters. Additionally, the intervals between the three outbursts of this repeater are quite long, and the features of each burst show significant variation. This may indicate that the FRB sources are unrelated and originate from different galaxies within the same direction or from the same galaxy. Furthermore, within Cat1, there are some localized non-repeaters, and we identified two of them as repeater candidates using both clustering algorithms. Continued observations of the host galaxies of these two FRBs may reveal additional repeating bursts in the future.
@@
We further analyzed the empirical relations that may exist within different clusters. For the clusters derived from k-means, we identified a significant ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ relation exclusive to cluster 2, as well as a $r-\gamma$ relation present only in clusters 0 and 1. For the clusters obtained through HDBSCAN, we found a notable ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation that exists solely in cluster 0, along with the $r-\gamma$ relation observed across all clusters. The spectral index $\gamma$ and spectral running $r$ are the shape parameters of FRB spectrum, described by a continuous power-law function (Pleunis et al., 2021; Planck Collaboration et al., 2020):
@@
where $I(\nu)$ is the intensity at spectral frequency $\nu$, A is the amplitude, and $\nu_{0}$ is the pivotal frequency, set at 400.1953125 MHz, the lower limit of the CHIME band. The strict $r-\gamma$ relation means that only one parameter can determine the morphology of FRB. We also noted that the $r-\gamma$ relation is similar in certain clusters of both non-repeaters and repeaters, such as clusters 0 and 1 from k-means and clusters 0 and 3 from HDBSCAN. This suggests that some non-repeaters and repeaters share comparable spectral characteristics. This finding aligns with observational evidence, particularly regarding FRB 20180910A, if it is indeed a genuine repeater.

### v2

@@ 5 Conclusion and Future Prospect
Machine learning is a powerful tool to classify FRBs. In this paper, we applied unsupervised learning methods, including dimensionality reduction and clustering algorithms, to differentiate between repeaters and non-repeaters in the first CHIME/FRB catalog (CHIME/FRB Collaboration et al., 2021) and the CHIME/FRB Collaboration (2023) catalog (CHIME/FRB Collaboration et al., 2023). We extracted 16 parameters from the FRBs to serve as input features for unsupervised learning, ensuring that the information from the FRBs was sufficiently comprehensive. Ultimately, we successfully identified several candidate repeaters among the non-repeaters. Using the UMAP+k-means method, we identified 269 non-repeaters as repeater candidates, with an estimated repeater source percentage of 61.7%. With the UMAP+HDBSCAN method, 141 non-repeaters were identified as repeater candidates, yielding a repeater source percentage of 37.9%. All repeater candidates are summarized in Appendix B. Additionally, we found that FRBs in repeater clusters and non-repeater clusters exhibit different distributions across several features, suggesting that repeaters and non-repeaters may belong to distinct categories.
@@
We used six FRB sources previously classified as non-repeaters but actually confirmed as repeaters to evaluate the predictive capability of our model. The UMAP + k-means method successfully predicted five of these sources, while the UMAP + HDBSCAN method successfully predicted four. The only exception was FRB 20180910A, which could not be predicted. The reason for this is that many of its characteristics, such as frequency bandwidth, spectral index, and spectral running, closely resemble those of non-repeaters, making it distinctly different from typical repeaters. Additionally, the intervals between the three outbursts of this repeater are quite long, and the features of each burst show significant variation. This may indicate that the FRB sources are unrelated and originate from different galaxies within the same direction or from different sources in the same galaxy. Furthermore, within Cat1, there are some localized non-repeaters, and we identified two of them as repeater candidates using both clustering algorithms. Continued observations of the host galaxies of these two FRBs may reveal additional repeating bursts in the future.
@@
We further analyzed the empirical relations that may exist within different clusters. For the clusters derived from k-means, we identified a significant (with $R^{2}>0.5$) ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ relation exclusive to cluster 2, as well as a $r-\gamma$ relation present only in clusters 0 and 1. For the clusters obtained through HDBSCAN, we found a notable ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation that exists solely in cluster 0, along with the $r-\gamma$ relation observed across all clusters. The spectral index $\gamma$ and the spectral running $r$ are the shape parameters of the FRB spectrum, described by a continuous power-law function (Pleunis et al., 2021; Planck Collaboration et al., 2020):
@@
where $I(\nu)$ is the intensity at spectral frequency $\nu$, A is the amplitude, and $\nu_{0}$ is the pivotal frequency, set at 400.1953125 MHz, the lower limit of the CHIME band. The strict $r-\gamma$ relation means that only one parameter can determine the morphology of FRB. Some clusters from k-means exhibit relatively weak but still noteworthy empirical relations, such as the ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ relation in clusters 0 and 1 and the $r-\gamma$ relation in cluster 2. Similarly, for HDBSCAN, the ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation is present but weak in cluster 1, while it is extremely weak in clusters 2, 3, and 4, suggesting that it is effectively nonexistent in the latter. These results suggest that the ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$ relation is more pronounced in repeater clusters, whereas the ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$ relation is more evident in non-repeater clusters. While the $r-\gamma$ relation is significant in both repeaters and non-repeaters, the noticeable differences in slopes and intercepts across different repeater clusters indicate that repeaters do not form a homogeneous group, and their properties may vary significantly.
We also applied the Chow test to assess statistical differences in the slopes and intercepts of empirical relationships across different clusters. For certain prominent empirical relationships, the Chow test $p$-values indicate that some repeater and non-repeater clusters share a common regression model, such as the $r-\gamma$ relation in clusters 0 and 1 from k-means and clusters 0 and 3 from HDBSCAN. This suggests that some repeaters and non-repeaters exhibit comparable spectral characteristics, a finding consistent with observational evidence—particularly in the case of FRB 20180910A, if it is indeed a genuine repeater. However, this also implies that if some repeaters behave similarly to non-repeaters in key spectral relationships, some FRBs currently classified as non-repeaters might actually be repeaters with undetected bursts due to observational limitations. Subsequently, when all repeater and non-repeater clusters are merged separately, the Chow test $p$-values reveal significantly distinct empirical relationships between the two groups, reinforcing the notion that FRBs can indeed be categorized into repeaters and non-repeaters.